OpenStudy (anonymous):

(1-cos^2x)(1+cot^2x)=1

OpenStudy (anonymous):

\[(1-\cos ^{2}x) (1+\cot ^{2}x)=1\] \[Formula : 1=\sin ^{2}x+\cos ^{2}x\] \[cosec ^{2}x=1+\cot ^{2}x\] \[(\sin ^{2}x+\cos ^{2}x-\cos ^{2}x)(cosec ^{2}x)=1 \] \[\sin ^{2}x \times cosec ^{2}x=1\] \[1=1\]

OpenStudy (anonymous):

@daya128 how about that ???

OpenStudy (anonymous):

wats cosec^2x?